# Chaos propagation in genetic algorithms: An optimal transport approach

### v1

@@ Abstract
Genetic algorithms are high-level heuristic optimization methods which enjoy great popularity thanks to their intuitive description, flexibility, and, of course, effectiveness. The optimization procedure is based on the evolution of possible solutions following three mechanisms: selection, mutation, and crossover. In this paper, we look at the algorithm as an interacting particle system and show that it is described by a Boltzmann-type equation in the many particles limit. Specifically, we prove a propagation of chaos result with a novel technique that leverages the optimal transport formulation of the Kantorovich–Rubinstein norm and naturally incorporates the crossover mechanism into the analysis. The convergence rate is sharp with respect to the number of particles.
keywords Genetic algorithms, interactive particle systems, Boltzmann equation, propagation of chaos, Wasserstein distance, Nanbu method
MSCcodes 82C22, 35Q20, 65C05, 90C59, 49Q22
@@ 1.1 Main result
converges to measure solutions of (1.2) with a rate. To quantify convergence, we consider the Kantorovich–Rubinstein norm [26] for signed Radon measures
$\|\mu\|_{\mathtt{KR}}:=\sup_{\textup{Lip}(\phi)\leq 1}\int\phi(x)\mu(\mathrm{d}x)\,\qquad\mu\in\mathcal{M}(\mathbb{R}^{d})$ (1.4)
which extends the $L^{1}$-Wasserstein distance between probability measures with finite first moments, $\mathbb{W}_{1}(f,g)=\|f-g\|_{\mathtt{KR}}$, $f,g\in\mathcal{P}_{1}(\mathbb{R}^{d})$ [33].
@@
$\mathbb{E}\|f^{N}_{0}-f_{0}\|_{\mathtt{KR}}\lesssim M_{q}^{1/q}(f_{0})\varepsilon_{1}(N)$ (1.5)
where $M_{q}(f_{0}):=\int|x|^{q}f_{0}(\mathrm{d}x)$, and $\varepsilon_{1}(N)\to 0$ as $N\to\infty$, is a rate which only depends on $N$. We refer to [16] or Section 2 for more details on this concentration inequality.
The main result of the paper is showing that the optimal convergence rate $\varepsilon_{1}(N)$ is maintained also at subsequent times, with an additional error of order $\mathcal{O}(\tau)$ given by the time discretization of the PDE. This means that particles generated by Algorithm 1 become independent on one another as $N\to\infty$, showing chaotic behavior.
@@
The fitness function is bounded, and satisfies $\mathbf{F}:\mathbb{R}^{d}\to[\underline{\mathbf{F}},\overline{\mathbf{F}}]$ with $\overline{\mathbf{F}}>\underline{\mathbf{F}}>0$. We denote by $\kappa_{\mathbf{F}}:=\max\mathbf{F}/\min\mathbf{F}$ the selection pressure.
@@
Let $f_{0}$ and $\mathbf{F}$ satisfy Assumptions 1.1 and 1.2, $\mu^{\gamma}=\textup{Unif}[0,1]^{d}$, and $\mu^{\xi}=\mathcal{N}(0,I_{d})$. Let $f\in C([0,T],\mathcal{P}_{q}(\mathbb{R}^{d}))$ be a weak measure solution to (1.2) in duality with $C_{b}(\mathbb{R}^{d})$ (see Definition 2.1) with initial data $f_{0}$, and time horizon $T>0$.
@@
$\sup_{t_{n}=n\tau\leq T}\mathbb{E}\|f^{N}_{n}-f(t_{n})\|_{\mathtt{KR}}\leq C(1+\kappa^{2}_{\mathbf{F}})Te^{C^{\prime}\kappa_{\mathbf{F}}T}\left(\,\varepsilon_{1}(N)\,+\,\tau\,\right)$
with $C=C(d,q,M_{q}(f_{0})^{1/q},\sigma),C^{\prime}=C^{\prime}(q)$ positive constants.
@@ 1.2 Literature review and novelty
The connection between interacting particle systems and genetic algorithms has been extensively studied in the series of works [10, 13, 12, 11]. GA is viewed as a numerical Monte Carlo approach to solve Feynman–Kac formulae [14] which is a class of more general dynamics also arising in nonlinear filtering methods and Sequential Monte Carlo samplers [8]. To obtain this interpretation, only the selection-mutation mechanism (without crossover) is considered. The dynamics is studied both in discrete and continuous time settings, see for example [13]. The authors address both stability with respect to the many particle limit $N\to\infty$, and long-time stability of the resulting measure-valued iteration. Such convergence analysis is typically based on martingales and semi-group techniques, the use of total-variation norms, and Zolotarev semi-norms [31, 9]. Oftentimes the fitness is assumed to be bounded as in Assumption (1.2), see for instance [13, Condition ($\mathcal{G}$), Eq. (22)]. The mechanism of crossover, where two selected particles are mixed to create a new one, is to our knowledge only considered in [11] by doubling the state space (in our setting, $\mathbb{R}^{d}\times\mathbb{R}^{d}$) in order not to break the Feynman–Kac-type description.
@@
We provide a concise proof of Theorem 1.3 by exploiting both the primal and dual formulation of the $L^{1}$-Wasserstein distance. With the first, we treat the crossover-mutation mechanism, thanks to the Lipschitzianity of the collision (1.1), while with the second we show stability of the selection procedure. The technique allows to optimally couple the parent particles at each step, similar to the technique employed in [17, 7], so that the optimal convergence rate of i.i.d. particles is not lost. The key argument is Lemma 3.1 which is a modified version of [4, Lemma 4.1] for weighted particles.
@@ 2 Notation and preliminaries
Random variables are assumed to be defined with respect to an underlying probability space $(\Omega,\mathcal{F},\mathbb{P})$. $X\sim f$ means that $X$ is a random variable with law $f$. For a given Borel set $A\subseteq\mathbb{R}^{d}$, $\textup{Unif}(A)$ is the uniform probability measure over $A$. For $\tau\in(0,1)$, we have $\textup{Bern}(\tau)=(1-\tau)\delta_{0}+\tau\delta_{1}$ with $\delta_{x}$ the Dirac delta measure centred at $x$. The indicator function for a given set $A$ is $\bm{1}_{A}$.
@@
With $\mathcal{P}(\mathbb{R}^{d})$ we denote the set of Borel probability measures over $\mathbb{R}^{d}$, and for $\phi:\mathbb{R}^{d}\to\mathbb{R}$ we write $\langle\phi,f\rangle=\int_{\mathbb{R}^{d}}\phi(x)f(\mathrm{d}x)$. Let $M_{p}(f):=\int|x|^{p}f(\mathrm{d}x),p\geq 1$ be the $p$-th moment of $f$, and $\mathcal{P}_{p}(\mathbb{R}^{d})$ be the subset of probabilities with finite $p$-th moments, which we equip with the $L^{p}$-Wasserstein distance [33]
$\mathbb{W}_{p}(f,g)=\left(\inf_{X\sim f,Y\sim g}\mathbb{E}|X-Y|^{p}\right)^{1/p}.$
@@
We write $f\in C([0,T],\mathcal{P}_{p}(\mathbb{R}^{d}))$ if the curve $f(t),t\in[0,T],$ is continuous with respect to $\mathbb{W}_{p}$. As mentioned in the Introduction, we have equivalence between the $L^{1}$-Wasserstein distance and the Kantorovich–Rubinstein norm, in the sense that
$\mathbb{W}_{1}(f,g)=\|f-g\|_{\mathtt{KR}}=\sup\left\{\langle\phi,f-g\rangle\,\middle|\,\textup{Lip}(\phi)\leq 1\right\}$ (2.1)
where $\textup{Lip}(\phi):=\sup_{x\neq y}|\phi(x)-\phi(y)|/|x-y|$. $C_{b}(\mathbb{R}^{d})$ is the set of bounded, continuous test functions, and $L^{1}(f)=\{\phi:\mathbb{R}^{d}\to\mathbb{R}\,|\,\langle|\phi|,f\rangle<\infty\}$.
We will sometimes denote $\{1,\dots,N\}$ with $[N]$, and $\Delta_{N}$ is the unit simplex $\Delta_{N}=\{\mathbf{w}\in\mathbb{R}^{N}\,|\,\sum_{i}w^{i}=1,w\geq 0\}$, with $\textup{int}(\Delta_{N})=\Delta_{N}\cap\mathbb{R}_{>0}^{N}$ the interior points where weights are strictly positive.
Definition 2.1 (Weak measure solution to (1.2)).
@@
Given $T\in(0,\infty)$, we say $f\in C([0,T],\mathcal{P}_{q}(\mathbb{R}^{d}))$ is a weak measure solution to (1.2) with initial data $f_{0}\in\mathcal{P}_{q}(\mathbb{R}^{d})$ if for any $\phi\in C_{b}(\mathbb{R}^{d})$ we have $\lim_{t\to 0^{+}}\langle\phi,f(t)\rangle=\langle\phi,f_{0}\rangle$ and
@@
We recall the empirical concentration inequality from [16, Theorem 1], in the specific case of $L^{1}$-Wasserstein distance, and under the assumption of having sufficient finite moments.
@@
$\mathbb{E}\mathbb{W}_{1}(f,f^{N})\leq CM_{q}^{1/q}(f)\varepsilon_{1}(N)$
@@
Let $f\in\mathcal{P}(\mathbb{R}^{d})$ and let $\mathbf{F}$ satisfy Assumption 1.2. Then for any $\phi\in L^{1}(f)$ with $\phi\geq 0$,
$\left\langle\phi,\frac{\mathbf{F}f}{\langle\mathbf{F},f\rangle}\right\rangle\leq\kappa_{\mathbf{F}}\langle\phi,f\rangle\,.$
@@
By Assumption 1.2 we have $0<\underline{\mathbf{F}}\leq\min\mathbf{F}\leq\max\mathbf{F}\leq\overline{\mathbf{F}}<+\infty$. Then for all $x$,
$0\leq\frac{\mathbf{F}(x)}{\langle\mathbf{F},f\rangle}\leq\frac{\max\mathbf{F}}{\min\mathbf{F}}=\kappa_{\mathbf{F}}\,$,
and, for any $\phi\in L^{1}(f)$ with $\phi\geq 0$,
$\left\langle\phi,\frac{\mathbf{F}f}{\langle\mathbf{F},f\rangle}\right\rangle=\int\phi(x)\frac{\mathbf{F}(x)}{\langle\mathbf{F},f\rangle}\,f(\mathrm{d}x)\leq\kappa_{\mathbf{F}}\int\phi(x)\,f(\mathrm{d}x)=\kappa_{\mathbf{F}}\langle\phi,f\rangle.$

### v2

@@ Abstract
Genetic algorithms are high-level heuristic optimization methods which enjoy great popularity thanks to their intuitive description, flexibility, and, of course, effectiveness. The optimization procedure is based on the evolution of possible solutions following three mechanisms: selection, mutation, and crossover. In this paper, we look at the algorithm as an interacting particle system and show that it is described by a Boltzmann-type equation in the many particles limit. Specifically, we prove a propagation of chaos result with a novel technique that leverages the optimal transport formulation of the bounded Lipschitz norm and naturally incorporates the crossover mechanism into the analysis. The convergence admits a rate with respect to the number of particles, corresponding to the optimal rate in the Wasserstein-1 distance.
keywords: Genetic algorithms, interactive particle systems, Boltzmann equation, propagation of chaos, Wasserstein distance, Nanbu method
MSCcodes: 82C22, 35Q20, 65C05, 90C59, 49Q22
@@ 1.1 Main result
converges to measure solutions of (1.2) with a rate. To quantify convergence, we consider the Bounded-Lipschitz (BL) norm [16] for signed Radon measures
$\|\mu\|_{\mathtt{BL}}:=\sup\left\{\int\phi(x)\mu(\mathrm{d}x)\,\middle|\,\textup{Lip}(\phi)\leq 1,\,\|\phi\|_{\infty}\leq\frac{1}{2}\right\}\,\qquad\mu\in\mathcal{M}(\mathbb{R}^{d})\,$, (1.4)
also known as flat metric between measures.
@@
$\mathbb{E}\|f^{N}_{0}-f_{0}\|_{\mathtt{BL}}\lesssim M_{q}^{1/q}(f_{0})\varepsilon_{1}(N)$ (1.5)
where $M_{q}(f_{0}):=\int|x|^{q}f_{0}(\mathrm{d}x)$, and $\varepsilon_{1}(N)\to 0$ as $N\to\infty$, is a rate which only depends on $N$. We refer to [19], or Lemma 2.2, for more details on this concentration inequality.
The main result of the paper is showing that the convergence rate $\varepsilon_{1}(N)$ is maintained also at subsequent times, with an additional error of order $\mathcal{O}(\tau)$ given by the time discretization of the PDE. This means that particles generated by Algorithm 1 become independent on one another as $N\to\infty$, showing chaotic behavior.
@@
The fitness function is globally Lipschitz continuous, bounded, and strictly positive. Specifically, it satisfies $\mathrm{Lip}(\mathbf{F}):=\sup_{x\neq y}|\mathbf{F}(x)-\mathbf{F}(y)|/|x-y|<+\infty$, and $\mathbf{F}:\mathbb{R}^{d}\to[\underline{\mathbf{F}},\overline{\mathbf{F}}]$ with $\overline{\mathbf{F}}>\underline{\mathbf{F}}>0$.
@@
Let $f_{0}$ and $\mathbf{F}$ satisfy Assumptions 1.1 and 1.2, $\mu^{\gamma}=\textup{Unif}[0,1]^{d}$, and $\mu^{\xi}=\mathcal{N}(0,I_{d})$. Let $f\in C([0,T],\mathcal{P}_{q}(\mathbb{R}^{d}))$ be a measure solution to (1.2) in duality with $C_{b}(\mathbb{R}^{d})$ (see Definition 2.1) with initial data $f_{0}$, and time horizon $T>0$.
@@
$\sup_{t_{n}=n\tau\leq T}\mathbb{E}\|f^{N}_{n}-f(t_{n})\|_{\mathtt{BL}}\leq CTe^{C^{\prime}T}\left(\,\varepsilon_{1}(N)\,+\,\tau\,\right)$
with $C=C(\mathrm{Lip}(\mathbf{F}),\underline{\mathbf{F}},\overline{\mathbf{F}},d,q,M_{q}(f_{0})^{1/q},\sigma),C^{\prime}=C^{\prime}(\mathrm{Lip}(\mathbf{F}),\underline{\mathbf{F}},\overline{\mathbf{F}},q)$ positive constants.
@@ 1.2 Literature review and novelty
The connection between interacting particle systems and genetic algorithms has been extensively studied in the series of works [11, 14, 13, 12]. GA is viewed as a numerical Monte Carlo approach to solve Feynman–Kac formulae [15] which is a class of more general dynamics also arising in nonlinear filtering methods and Sequential Monte Carlo samplers [9]. To obtain this interpretation, only the selection-mutation mechanism (without crossover) is considered. The dynamics is studied both in discrete and continuous time settings, see for example [14]. The authors address both stability with respect to the many particle limit $N\to\infty$, and long-time stability of the resulting measure-valued iteration. Such convergence analysis is typically based on martingales and semi-group techniques, the use of Total Variation norm, and Zolotarev semi-norms [37, 10]. Oftentimes the fitness is assumed to be bounded as in Assumption (1.2), see for instance [14, Condition ($\mathcal{G}$), Eq. (22)]. The mechanism of crossover, where two selected particles are mixed to create a new one, is to our knowledge only considered in [12] by doubling the state space (in our setting, $\mathbb{R}^{d}\times\mathbb{R}^{d}$) in order not to break the Feynman–Kac-type description.
@@
We provide a concise proof of Theorem 1.3 by exploiting both the primal and dual formulation of a $L^{1}$-Wasserstein distance $\mathbb{W}_{1}^{\wedge}$ with cost given by
${D^{\wedge}}(x,y):=|x-y|\wedge 1=\min\{|x-y|,1\}\,$, (1.6)
which coincides with the BL metric, $\mathbb{W}_{1}^{\wedge}(f,g):=\|f-g\|_{\mathtt{BL}}$, between two probability measures $f,g$, see Section 2 for more details. The distance $D^{\wedge}$} sits in between |xy||x-y| and $\bm{1}_{x\neq y}=\delta(x-y)$, leading to a metric which inherits benefits from both the usual Wasserstein-1 metric, and the Total Variation norm. With the optimal transport formulation, we treat the crossover-mutation mechanism, thanks to the Lipschitzianity of the collision (1.1), while with the dual one we show stability of the selection procedure. The technique allows to optimally couple the parent particles at each step, similar to the technique employed in [20, 8], so that the convergence rate of i.i.d. particles is not lost. The key argument is Lemma 3.1 which is a modified version of [4, Lemma 4.1] for weighted particles. The distance (1.6) has been used in [23, 18, 5], while generalizations beyond probability measures have been considered in [35, 34].
@@ 2 Notation and preliminaries
Random variables are assumed to be defined with respect to an underlying probability space $(\Omega,\mathcal{F},\mathbb{P})$. $X\sim f$ means that $X$ is a random variable with law $f$. For a given Borel set $A\subseteq\mathbb{R}^{d}$, $\textup{Unif}(A)$ is the uniform probability measure over $A$. For $\tau\in(0,1)$, we have $\textup{Bern}(\tau)=(1-\tau)\delta_{0}+\tau\delta_{1}$ with $\delta_{x}$ the Dirac delta measure centred at $x$. The indicator function for a given set $A$ is $\bm{1}_{A}$. We will sometimes denote $\{1,\dots,N\}$ with $[N]$, and $\Delta_{N}$ is the unit simplex $\Delta_{N}=\{\mathbf{w}\in\mathbb{R}^{N}\,|\,\sum_{i}w^{i}=1,w\geq 0\}$, with $\textup{int}(\Delta_{N})=\Delta_{N}\cap\mathbb{R}_{>0}^{N}$ the interior points where weights are strictly positive. We use $a\wedge b:=\min(a,b)$, and $a\vee b=\max\{a,b\}$ for the minimum and maximum operations.
2.1 Wasserstein distances
@@
With $\mathcal{P}(\mathbb{R}^{d})$ we denote the set of Borel probability measures over $\mathbb{R}^{d}$, and for $\phi:\mathbb{R}^{d}\to\mathbb{R}$ we write $\langle\phi,f\rangle=\int_{\mathbb{R}^{d}}\phi(x)f(\mathrm{d}x)$. Let $M_{p}(f):=\int|x|^{p}f(\mathrm{d}x),p\geq 1$ be the $p$-th moment of $f$, and $\mathcal{P}_{p}(\mathbb{R}^{d})$ be the subset of probabilities with finite $p$-th moments. For a given cost function $c$, and $p\geq 1$, we consider the Wasserstein distances [39]
$\mathbb{W}_{p}^{c}(f,g)=\left(\inf_{X\sim f,Y\sim g}\mathbb{E}\left[c(X,Y)^{p}\right]\right)^{1/p}.$
@@
We write $f\in C([0,T],\mathcal{P}_{p}(\mathbb{R}^{d}))$ if the curve $f(t),t\in[0,T],$ is continuous with respect to $\mathbb{W}_{p}^{|\cdot|}$ (the usual $L^{p}$-Wasserstein distance).
Let $\textup{Lip}^{d}(\phi):=\sup_{x\neq y}|\phi(x)-\phi(y)|/d(x,y)$ denote the Lipschitz constant of a function $\phi$ with respect to a distance $d$. By Kantorovich–Rubinstein duality [39, Theorem 1.14], when the cost $c$ is a distance, it holds
$\mathbb{W}_{1}^{c}(f,g)=\sup\left\{\langle\phi,f-g\rangle\,\middle|\,\textup{Lip}^{c}(\phi)\leq 1\right\}\,.$
If c(x,y)=|xy|c(x,y)=|x-y| one gets that $\mathbb{W}_{1}^{|\cdot|}$ corresponds to the Kantorovich–Rubinstein (KR) norm (duality with $\mathrm{Lip}(\phi):=\mathrm{Lip}^{|\cdot|}(\phi)\leq 1$), while if $c(x,y)=\delta(x-y)=\bm{1}_{x\neq y}$, $\mathbb{W}^{\delta}_{1}$ corresponds the Total-Variation (TV) norm (duality with $\|\phi\|_{\infty}:=\sup_{x}\phi(x)\leq 1$), see, for instance, [36, Section 3].
For the distance function (1.6) of our interest, ${D^{\wedge}}(x,y)=|x-y|\wedge 1$, condition $\mathrm{Lip}^{D^{\wedge}}(\phi)\leq 1$ leads to
$|\phi(x)-\phi(y)|\leq{D^{\wedge}}(x,y)=|x-y|\wedge 1\,\qquad\textup{for all}\quad x,y\in\mathbb{R}^{d}\,.$
Therefore, a test function $\phi$ needs to satisfy $\mathrm{Lip}(\phi)=\mathrm{Lip}^{|\cdot|}(\phi)\leq 1$, and $\textup{osc}(\phi):=\sup\phi-\inf\phi\leq 1$. Since $\langle\phi,f-g\rangle$ is invariant under translations, condition $\textup{osc}(\phi)\leq 1$ is equivalent to $\|\phi\|_{\infty}\leq 1/2$, as in the definition of the BL norm (1.4). Therefore, we obtain that for
$\mathbb{W}_{1}^{\wedge}(f,g):=\mathbb{W}^{D^{\wedge}}_{1}(f,g)=\inf_{X\sim f,Y\sim g}\mathbb{E}\left[|X-Y|\wedge 1\right]$
it holds
$\mathbb{W}_{1}^{\wedge}(f,g)=\|f-g\|_{\mathtt{BL}}\,.$
Moreover, since ${D^{\wedge}}(x,y)\leq|x-y|$, we also have $\mathbb{W}_{1}^{\wedge}\leq\mathbb{W}_{1}^{|\cdot|}$. Generalisations to this optimal transport formulation of the BL norm have been studied in [35, 34].
2.2 Auxiliary lemmas
We collect some definitions and auxiliary results that will be used throughout the paper.
Definition 2.1 (Measure solution to (1.2)).
@@ 2.2 Auxiliary lemmas
Given $T\in(0,\infty)$, we say $f\in C([0,T],\mathcal{P}_{q}(\mathbb{R}^{d}))$ is a measure solution to (1.2) with initial data $f_{0}\in\mathcal{P}_{q}(\mathbb{R}^{d})$ if for any $\phi\in C_{b}(\mathbb{R}^{d})$ we have $\lim_{t\to 0^{+}}\langle\phi,f(t)\rangle=\langle\phi,f_{0}\rangle$ and
@@
We recall the empirical concentration inequality from [19, Theorem 1], in the specific case of Wasserstein distance $\mathbb{W}^{|\cdot|}_{1}$, and under the assumption of having sufficient finite moments. Note that since $\mathbb{W}_{1}^{\wedge}\leq\mathbb{W}_{1}^{|\cdot|}$, the same inequalities holds of the BL norm as well.
@@
$\mathbb{E}\mathbb{W}^{|\cdot|}_{1}(f,f^{N})\leq CM_{q}^{1/q}(f)\varepsilon_{1}(N)$
@@
Lemma 2.3 (Crossover-mutation stability).
For any fixed $\gamma\in[0,1]^{d}$ and $\sigma>0,\xi\in\mathbb{R}^{d}$, the crossover-mutation update (1.1) is Lipschitz in $x,x_{*}$ with respect to the metric $D^{\wedge}$}.
Proof.
Since $\gamma\in[0,1]^{d}$, we have $|\gamma\odot z|\leq|z|$ for any $z\in\mathbb{R}^{d}$. Then, consider $x^{\prime}$ and $y^{\prime}$ defined as in (1.1), we have
$|x^{\prime}-y^{\prime}|$ $=\left|(1-\gamma)\odot x+\gamma\odot x_{*}+\xi\sigma-\Big((1-\gamma)\odot y+\gamma\odot y_{*}+\xi\sigma\Big)\right|$
$=|(1-\gamma)\odot(x-y)+\gamma\odot(x_{*}-y_{*})|\leq|x-y|+|x_{*}-y_{*}|\,.$
Then, Lipschitzianity with respect to ${D^{\wedge}}(x,y)=|x-y|\wedge 1$ follows by
$|x^{\prime}-y^{\prime}|\wedge 1\leq(|x-y|+|x_{*}-y_{*}|)\wedge 1\leq|x-y|\wedge 1+|x_{*}-y_{*}|\wedge 1\,.$
@@
Let $\mathbf{F}$ satisfy Assumption 1.2. There exists a positive constant $C_{\mathbf{F}}=C_{\mathbf{F}}(\underline{\mathbf{F}},\overline{\mathbf{F}},\mathrm{Lip}(\mathbf{F}))$ such that
$\mathbb{W}_{1}^{\wedge}\left(\frac{\mathbf{F}f}{\langle\mathbf{F},f\rangle},\frac{\mathbf{F}g}{\langle\mathbf{F},g\rangle}\right)\leq C_{\mathbf{F}}\mathbb{W}_{1}^{\wedge}(f,g)\quad\textup{for any}\quad f,g\in\mathcal{P}(\mathbb{R}^{d})\,.$
@@
We use the dual formulation (1.4) of $\mathbb{W}_{1}^{\wedge}$, and so consider a test function $\phi$ with $\mathrm{Lip}(\phi)\leq 1$ and $\|\phi\|_{\infty}\leq 1/2$. By triangular inequality we have
$\left\langle\phi,\frac{\mathbf{F}f}{\langle\mathbf{F},f\rangle}-\frac{\mathbf{F}g}{\langle\mathbf{F},g\rangle}\right\rangle\leq\left|\left\langle\phi,\frac{\mathbf{F}f}{\langle\mathbf{F},f\rangle}-\frac{\mathbf{F}g}{\langle\mathbf{F},f\rangle}\right\rangle\right|+\left|\left\langle\phi,\frac{\mathbf{F}g}{\langle\mathbf{F},f\rangle}-\frac{\mathbf{F}g}{\langle\mathbf{F},g\rangle}\right\rangle\right|=:I_{1}+I_{2}\,.$
Since dividing $\phi\mathbf{F}$ by $\mathrm{Lip}(\phi\mathbf{F})\vee 2\|\phi\mathbf{F}\|_{\infty}$ turns it into a test function for the BL norm, it holds
$\langle\phi\mathbf{F},f-g\rangle\leq\left(\mathrm{Lip}(\phi\mathbf{F})\vee 2\|\phi\mathbf{F}\|_{\infty}\right)\|f-g\|_{\mathtt{BL}}\,.$
Given that $\|\phi\mathbf{F}\|_{\infty}\leq\overline{\mathbf{F}}/2$ and $\mathrm{Lip}(\phi\mathbf{F})\leq\mathrm{Lip}(\mathbf{F})/2+\overline{\mathbf{F}}$, we get the following bound for $I_{1}$
$I_{1}\leq\frac{1}{\underline{\mathbf{F}}}|\langle\phi,\mathbf{F}f-\mathbf{F}g\rangle|\leq\frac{\mathrm{Lip}(\phi\mathbf{F})\vee 2\|\phi\mathbf{F}\|_{\infty}}{\underline{\mathbf{F}}}\|f-g\|_{\mathtt{BL}}\leq\frac{(\mathrm{Lip}(\mathbf{F})/2+\overline{\mathbf{F}})}{\underline{\mathbf{F}}}\|f-g\|_{\mathtt{BL}}\,$,
where we also used $\langle\mathbf{F},f\rangle\geq\underline{\mathbf{F}}$ in the first inequality. Using the same type of estimates, we have for $I_{2}$
$I_{2}$ $=\left|\left\langle\phi,\frac{\mathbf{F}g}{\langle\mathbf{F},f\rangle}-\frac{\mathbf{F}g}{\langle\mathbf{F},g\rangle}\right\rangle\right|$ $=\left(\frac{1}{\langle\mathbf{F},f\rangle}-\frac{1}{\langle\mathbf{F},g\rangle}\right)\left|\left\langle\phi,\mathbf{F}g\right\rangle\right|$
$\leq\frac{|\langle\mathbf{F},g\rangle-\langle\mathbf{F},f\rangle|}{\langle\mathbf{F},f\rangle\langle\mathbf{F},g\rangle}\|\phi\mathbf{F}\|_{\infty}$ $\leq\frac{\overline{\mathbf{F}}(\mathrm{Lip}(\mathbf{F})\vee 2\overline{\mathbf{F}})}{2\underline{\mathbf{F}}^{2}}\|f-g\|_{\mathtt{BL}}\,.$
By collecting the two bounds, and simplifying slightly, one obtains the desired estimate with
$C_{\mathbf{F}}:=\frac{1}{\underline{\mathbf{F}}}\left(1+\frac{\overline{\mathbf{F}}}{\underline{\mathbf{F}}}\right)(\mathrm{Lip}(\mathbf{F})+2\overline{\mathbf{F}})\,.$